A standard hard water (SHW) sample contains 0.50 mg of CaCO3 per mL. 50 mL of SHW consumed 45 mL of EDTA using EBT indicator at end point. 25 mL of sample water consumed 15 mL of EDTA, and after boiling and filtering, 25 mL of this boiled water sample consumed 4 mL of EDTA. Calculate the total, permanent and temporary hardness of water.

Grade:12th pass

1 Answers

Rahul
4 years ago
1ml of Standard hard water = 0.5mg of CaCO3 50ml of Standard hard water = 25 mg of CaCO3 Also, 50 ml of standard hard water = 45ml of EDTA this means1 ml EDTA = 25/45 mg CaCO3 --------------(1)Also, 25 ml of sample water = 15 ml EDTA Using (1)25ml of sample water = 25/45 * 15mg CaCO3 1000ml of sample water = 1000/25 * 25/45 *15 mg CaCO3 = 1000/3 mg CaCO3 Total Hardness = 1000/3 mg CaCO3 Now, 25ml of boiled water sample = 4ml of EDTA solutionTherefore, 1000ml of boiled water sample = 4/25 *25/45 * 1000 = 800/9 mg CaCO3 So, Permanent Hardness = 800/9 mg CaCO3. As we know, Total Hardness = Permanent Hardness + Temporary Hardness So Temporary Hardness = 1000/3 - 800/9 =2200/9 So, Total hardness =333.33 mg CaCO3 Permanent Hardness = 88.88 mg CaCO3 Temporary Hardness = 244.44 mg CaCO3 Hope it helps For further queries call